OpenStudy (anonymous):

(1/3+squarerootof7/6 i)^2 It wants it in standard form. Could anyone assist me, I've been doing hw all day and I cant think anymore.

OpenStudy (anonymous):

Wait\[(\frac{ 1 }{ 3 }+i \sqrt{\frac{ 7 }{ 6 }})^2\] Is it this?

OpenStudy (anonymous):

There's so many forms of what you just wrote. You need to put brackets so I know the correct form.

OpenStudy (anonymous):

Could you "draw" it?

OpenStudy (anonymous):

My mistake. [(1/3)+(square root of 7/6)i]^2 Write the expression in standard form a+bi

OpenStudy (anonymous):

And Standard form means this: \[x+iy\] or \[a+ib\] where x and a are real and y and b are imaginary.

OpenStudy (anonymous):